OpenStudy (anonymous):

How would you solve 9x^2+9x+2=0 by factoring? I am stuck :(

OpenStudy (campbell_st):

well do it this way... works every time for a quadratic in the form \[ax^2 + bx + c \] multiply a and c so in your question 9 x 2 = 18 now find the factors of 18 that add to b... or in your question 9 then the solution is (ax + factor 1)(ax + factor 2) -------------------------- a in you question its (9x + factor 1)(9x + factor 2) -------------------------- 9 you'll find that each binomial will have common factors that cancel with the denominator

OpenStudy (campbell_st):

so 1st find the factors

OpenStudy (dangerousjesse):

Solve for x over the real numbers: \(9 x^2+9 x+2 = 0\) Factor the left hand side. The left hand side factors into a product with two terms: \((3 x+1) (3 x+2) = 0\) Solve each term in the product separately. Split into two equations: \(3 x+1 = 0\) or \(3 x+2 = 0\) Look at the first equation: Isolate terms with x to the left hand side. Subtract 1 from both sides: \(3 x = -1\) or \(3 x+2 = 0\) Solve for x. Divide both sides by 3: \(x = \frac{-1}{3}\) or \(3 x+2 = 0\) Look at the second equation: Isolate terms with x to the left hand side. Subtract 2 from both sides: \(x = \frac{-1}{3}\) or \(3 x = -2\) Solve for x. Divide both sides by 3...

OpenStudy (anonymous):

LOL, I'm a real idiot. I was wondering how any factors of 2 could be 9 but I just realized that there are 3x and not x. Thanks to both of you

OpenStudy (anonymous):

I didn't really get campbell_st's explanation

OpenStudy (campbell_st):

well to finish it... if you multiply a and c its 9 x 2 = 18 factors of 18 that add to 9... the coefficient of b are 6 and 3 so the next step is (9x + 6)(9x + 3) --------------- = 0 9 remove the common factors from each binomial 3(3x + 2)3(3x + 1) ------------------- = 0 9 cancel the common factors and you get (3x + 2)(3x + 1) = 0 this is a really effective method for dealing with any non-monic quadratic that can be factored.... it beats guess and check perhaps you should show your teacher...

OpenStudy (anonymous):

Ok, I'll write it down!!! :)

OpenStudy (campbell_st):

tell you teacher it works every time

OpenStudy (anonymous):

I'll do that too

OpenStudy (anonymous):

I'm kind of torn on who to give the medal

OpenStudy (anonymous):

You both did a good job

OpenStudy (anonymous):

I chose DangerousJesse because I understood his explanation, but I don't know now